OpenStudy (anonymous):

Here we go.....6(e^x/e^x-2)^-1/2

OpenStudy (anonymous):

\[6\left( \frac{ e ^{x} }{ e ^{x} -2} \right)^{-\frac{ 1 }{ 2 }} \text{ like this?}\]

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

you can start simplifying by flipping the terms inside the parenthteses and seeing what you get